OpenStudy (anonymous):

Help, please! Solve the equation (4x + 1)(3x - 2) = 0 A.{-4, 1.5} B.{-1/4, 2/3} C.{1/4, 2/3} D.{-4, 3}

OpenStudy (precal):

4x+1=0 -1 -1 4x=-1 divide 4 to both sides x=-1/4

OpenStudy (precal):

Now solve 3x-2=0 and you are done

OpenStudy (anonymous):

When multiple factors are multiplied to equal 0, you can solve for the variable by setting each factor equal to zero, then solving.

OpenStudy (anonymous):

B.{-1/4, 2/3}